1.    A particle of mass m moves in the x y-plane. Its position vector as a function of time is given by  where ω = angular velocity of the Find Velocity



1
Expert's answer
2021-06-14T15:01:31-0400

Gives

Mass (m)=m

r(t)=xi+yj

"v=r\times w"

"w=w\hat z"

We know that

"v=\begin{vmatrix}
 \hat i & \hat j&\hat k \\
 x&y&0 \\
0&0&w_z\\

\end{vmatrix}"

Solve it


"v=\hat i(w_zy-0)-\hat j(xw_z-0)+\hat k(0-0)"

"v=\hat iw_zy-\hat jxw_z"